# Role of sub-barrier dynamics in coherent and incoherent radiation during tunneling ionization

### v1

@@ IV Estimation of radiation intensities
where $w_{i}\sim(\kappa^{2}/2)(E_{a}/E_{0})\exp[-2E_{a}/(3E_{0})]$ is the laser ionization rate, $T=2\pi/\omega_{0}$ is the laser period. The probability of the coherent Brunel photon emission per laser period with the energy $\omega$ is
@@
For $E_{0}=0.1$ (the laser intensity $3.5\times 10^{14}$ W/cm2), $E_{0}/E_{a}\sim 0.1$, and $a_{0}\sim 1.5\times 10^{-2}$, $\frac{I_{p}}{\omega_{0}}\sim 10$, $e^{-\frac{2E_{a}}{3E_{0}}}\sim 1.3\times 10^{-3}$, $w_{i}T\sim 10^{-2}$, we obtain $w_{\rm sp}/w_{\rm coh}\sim 10^{2}$, which is roughly in accordance with the spectra of Figs. 1 and 4.
@@
V Conclusion
We have investigated the coherent and incoherent radiation during tunneling ionization. The coherent radiation in the near-zero-frequency ($\omega\sim 0.2$ eV) domain is dominated by the continuum dynamics via the $j_{11}$ term in the current. We have shown that due to the under-the-barrier motion the radiation can be enhanced. The nonadiabaticity increases the effect of the under-the-barrier dynamics at large Keldysh parameters, however generally decreasing the radiation yield.
The spontaneous radiation is not affected by the asymmetry of the driving laser pulse. The characteristic width of the spontaneous radiation is determined by the Keldysh time $\Delta\omega\sim 1/\tau_{K}$. There are two contributions into the spontaneous radiation: due to the direct transition during tunneling (with the amplitude $m_{0}$) investigated in Ref [14], and due to additional motion in the continuum (with the amplitude $m_{1}$). At larger frequencies $\omega\gtrsim 1/\tau_{K}$, $m_{0}$ and $m_{1}$ terms compensate each other and the total radiation is suppressed. At low frequencies $\omega\lesssim 1/\tau_{K}$, the radiation is dominated by the continuum dynamics via the amplitude $m_{1}$. Thus, the contribution of the direct transition in the radiation, corresponding to the result of Ref. [14], essentially describes the spontaneous radiation only near the cutoff.

### v2

@@ V Estimation of radiation intensities
where $w_{i}\sim(\kappa^{2}/2)(E_{0}/E_{a})\exp[-2E_{a}/(3E_{0})]$ is the laser ionization rate, and $T=2\pi/\omega_{0}$ the laser period. The probability of the coherent Brunel photon emission per laser period with the energy $\omega$ is
@@
For $E_{0}=0.1$ (the laser intensity $3.5\times 10^{14}$ W/cm2), $E_{0}/E_{a}\sim 0.1$, and $a_{0}\sim 1.5\times 10^{-2}$, $\frac{I_{p}}{\omega_{0}}\sim 10$, $e^{-\frac{2E_{a}}{3E_{0}}}\sim 1.3\times 10^{-3}$, $w_{i}T\sim 10^{-2}$, we obtain $w_{\rm sp}/w_{\rm coh}\sim 10^{2}$. The latter ratio is roughly in accordance with spectra in Figs. 1 and 3.
@@
VI Conclusion
We have investigated the coherent and incoherent radiation during tunneling ionization. The coherent radiation in the near-zero-frequency THz domain ($\omega\sim 0.2$ eV) is dominated by the continuum dynamics via the quantum free-free transition $J_{11}$ term in the current. We have shown that due to the quantum dynamics the radiation is modified especially in the adiabatic limit.
The spontaneous radiation is not affected by the asymmetry of the driving laser pulse. The characteristic width of the spontaneous radiation is determined by the Keldysh time $\Delta\omega\sim 1/\tau_{K}$. There are two contributions into the spontaneous radiation: due to the direct transition during tunneling (with the amplitude $m_{0}$) investigated in Ref [30], and due to additional motion in the continuum (with the amplitude $m_{1}$). At larger frequencies $\omega\gtrsim 1/\tau_{K}$, $m_{0}$ and $m_{1}$ terms compensate each other and the total radiation is suppressed. At low frequencies $\omega\lesssim 1/\tau_{K}$, the radiation is dominated by the continuum dynamics via the amplitude $m_{1}$. Thus, the contribution of the direct transition in the radiation essentially describes the spontaneous radiation only near the cutoff.